# Sampling and estimation on manifolds using the Langevin diffusion

### v2

@@ 6.2. Sampling on the manifold of SPD matrices
We consider the non-compact manifold of symmetric positive definite (SPD) matrices, and demonstrate numerically that the proposed Riemannian Langevin algorithm results in the desired error bounds when sampling from two distributions which satisfy the assumptions discussed in Section 5. In contrast to the unit sphere in Section 6.1 the algorithm is implemented using a closed form expression for the exponential map.
Consider the general linear group $GL(m)$ consisting of the set of $m\times m$ real invertible matrices equipped with the group operation of matrix multiplication. Denote by $M=\mathcal{P}_{m}$, the manifold of $m\times m$ symmetric positive definite matrices $\mathcal{P}_{m}=\{X\in GL(m):X^{\intercal}=X,\;y^{\intercal}Xy>0,\;y\in\mathbb{R}^{m}\}$, an open cone of dimension $q=m(m+1)/2$.
@@
Affine invariance of the metric ensures that the Riemannian distance $\rho$ and volume form $\mathrm{dvol}_{g}$ will also be invariant under affine transformations. The metric $g$ gives $\mathcal{P}_{m}$ a Riemannian structure with non-positive sectional curvature, and $\mathcal{P}_{m}$ is a non-compact, geodesically complete metric space by the Hopf–Rinow theorem (e.g., Jost, 2008). Moreover, by the Cartan–Hadamard theorem (e.g., Kobayashi and Nomizu, 1969), the inverse exponential map is globally defined, and the cut locus of every point is empty, i.e., every two points in $\mathcal{P}_{m}$ have a unique geodesic connecting them. This means that only a single chart is needed to cover $\mathcal{P}_{m}$.
@@
$\rho(X_{1},X_{2})=\sqrt{\sum_{i=1}^{n}(\log r_{i})^{2}},$
@@
6.2.1. Riemannian–Gaussian distribution
The Riemannian–Gaussian distribution can be viewed as the generalisation of the isotropic normal distribution on $\mathbb{R}^{q}$ to $\mathcal{P}_{m}$, where the Euclidean distance is replaced by the Riemannian distance in the probability density function:
@@ 6.2.1. Riemannian–Gaussian distribution
The Riemannian–Gaussian distribution satisfies Assumptions Assumption and Assumption, as seen in Example 5.2 (i). This ensures that the Langevin diffusion converges to its invariant measure exponentially fast. From (6.8) the gradient of $\phi$ is
@@
Algorithm 6.2 Algorithm to sample from Riemannian–Gaussian distribution on $\mathcal{P}_{3}$ with $O=I_{3}$, $\sigma=\frac{1}{\sqrt{2}}$ and $X_{0}=\mathsf{hvec}^{-1}((2,4,2,1,1,0)^{\intercal})$.
@@
Figure 2. Sampling from the Riemannian–-Gaussian distribution on $\mathcal{P}_{3}$: log-log plot of the estimation error err against $h$. Error bars correspond to Monte Carlo error MCerr. The reference red line has gradient 1. The parameters for the Riemannian–Gaussian are as in Table 3.
@@ 6.2.2. Distribution with non-convex potential
with the normalisation constant $Z_{m}$. This distribution is referred to as one with a double-well potential. Similar to the Riemannian–Gaussian distribution, the constant $Z_{m}$ is obtained by evaluating
@@ 7. Discussion and concluding remarks
The main focus is on sampling from, and computing expectation of functions of, a target distribution. As a consequence, we demonstrate how the concept of weak approximations for SDEs in $\mathbb{R}^{q}$ and the flat torus $\mathbb{T}^{q}$, and the corresponding proof techniques in deriving weak error bounds, can be transferred to the case of compact Riemannian manifolds via the covariant Taylor expansion. The PDE-based proof technique enables obtaining weak error bounds that match the optimal bounds in $\mathbb{R}^{q}$ and flat manifolds; these are superior to the weak error bound currently available in the literature, which follow by first deriving strong approximation (e.g., convergence in expected squared-distance) error bounds. The price we pay for our approach is in the absence of explicit dimension-dependent constants and the spectral gap in the derived bounds. The Talay-Tubaro extrapolation in Section 4.4 may be used to extract more information on the constants, and this constitutes ongoing work.
@@ Acknowledgements
The authors were supported by EPSRC grant no. EP/X022617/1. KB acknowledges support from grants EPSRC EP/V048104/1, NSF 2015374 and NIH R37-CA21495.
Appendix A Proofs
Appendix B
@@ Appendix A Proofs
(B.1) $F(s)=F(0)+s\frac{\text{d}}{\text{d}s}F(0)+\frac{s^{2}}{2}\frac{\text{d}^{2}}{\text{d}s^{2}}F(0)+\frac{s^{3}}{6}\frac{\text{d}^{3}}{\text{d}s^{3}}F(0)+\frac{s^{4}}{24}\frac{\text{d}^{4}}{\text{d}s^{4}}F(\alpha),$
@@
(B.2) $\dfrac{\text{d}}{\text{d}s}F(s)=\dfrac{\text{d}}{\text{d}s}f(c(s))=\text{d}f(c(s))(\dot{c}(s)).$
@@
(B.4) $\;\;\;+\text{d}f(c(s))(D_{\dot{c}(s)}D_{\dot{c}(s)}\dot{c}(s)),$
where we have utilized the fact that second covariant derivative of functions is symmetric $(0,2)-$tensor. Evaluating (B.2), (B), (B) at $s=0$ and substituting them into (B.1), we obtain the following Taylor formula of $f$ along the curve $c$:
@@
(B.5) $\;\;\;+\dfrac{s^{4}}{24}D^{4}f(c(\alpha))(\dot{c}(\alpha),\dot{c}(\alpha),\dot{c}(\alpha),\dot{c}(\alpha)),$
@@
Appendix C Proofs of Theorems
@@
Lemma C.1 (One-step error).
@@
(C.3) $u(t+h,c(s))=u(t,c(s))+h\frac{\partial u}{\partial t}(t,c(s))+\frac{h^{2}}{2}\frac{\partial^{2}u}{\partial t^{2}}(t(\alpha_{1}),c(s)),$
@@
Applying the covariant Taylor formula (B) to the function $u(t,c(s))$ along the curve $c$, we get
@@
(C.4) $+\dfrac{s^{4}}{24}D^{4}u(t,c(\alpha_{2}))(\dot{c}(\alpha_{2}),\dot{c}(\alpha_{2}),\dot{c}(\alpha_{2}),\dot{c}(\alpha_{2})),$
@@
(C.5) $\;\;\;+\frac{s^{2}}{2}\bigg{(}D^{2}\frac{\partial u}{\partial t}(t,c(\alpha_{3}))(\dot{c}(\alpha_{3}),\dot{c}(\alpha_{3}))+\text{d}\dfrac{\partial u}{\partial t}(t,c(\alpha_{3}))(D_{\dot{c}(\alpha_{3})}\dot{c}(\alpha_{3}))\bigg{)},$
@@
$+\frac{s^{4}}{24}D^{4}u(t,c(\alpha_{2}))(\dot{c}(\alpha_{2}),\dot{c}(\alpha_{2}),\dot{c}(\alpha_{2}),\dot{c}(\alpha_{2}))$
@@
(C.6) $+\frac{h^{2}}{2}\frac{\partial^{2}u}{\partial t^{2}}(t(\alpha_{1}),c(s)).$
@@
$+\frac{h^{2}}{24}D^{4}u(t,c(\alpha_{2}))(\dot{c}(\alpha_{2}),\dot{c}(\alpha_{2}),\dot{c}(\alpha_{2}),\dot{c}(\alpha_{2}))$
@@
(C.7) $+\frac{h^{2}}{2}\frac{\partial^{2}u}{\partial t^{2}}(t(\alpha_{1}),c(s))\bigg{]}.$
@@
(C.8) $\mathbb{E}[\xi^{i}]=0,\;\mathbb{E}[\xi^{i}\xi^{j}]=\delta_{ij},\;\mathbb{E}[\xi^{i}\xi^{j}\xi^{k}]=0,\;\mathbb{E}[(\xi^{i})^{2}(\xi^{j})^{2}]=1\text{ for }i\neq j,$
@@
(C.9) $=-\frac{h^{3/2}}{8}(\nabla\phi)^{i}(\nabla\phi)^{j}(\nabla\phi)^{k}-\frac{3h^{1/2}}{2}(\nabla\phi)^{i}g^{jk},$
@@
(C.10) $\mathbb{E}(h^{1/2}Du(V))$ $=-\frac{h}{2}g(\nabla\phi,\nabla u),$
@@
(C.11) $=h\Delta_{M}u+\frac{h^{2}}{4}D^{2}u(\nabla\phi,\nabla\phi),$
@@
(C.12) $\;\;\;\;-\frac{3h^{2}}{2}D^{3}u(\partial_{i},\partial_{j},\partial_{k})(\nabla\phi)^{i}g^{jk},$
(C.13) $\mathbb{E}\Big{(}h^{3/2}D\frac{\partial u}{\partial t}(V)\Big{)}$ $=h^{3/2}D\frac{\partial u}{\partial t}(\partial_{i})\mathbb{E}(V^{i})=\frac{h^{2}}{2}D\frac{\partial u}{\partial t}(\partial_{i})(\nabla\phi)^{i}.$
Substituting (C.10)-(C.13) in (C.7), we get
@@
$\;\;\;\;+\mathbb{E}\bigg{(}\frac{h^{2}}{24}D^{4}u(t,c(\alpha_{2}))(\dot{c}(\alpha_{2}),\dot{c}(\alpha_{2}),\dot{c}(\alpha_{2}),\dot{c}(\alpha_{2}))$
@@
(C.14) $\;\;\;\;+\frac{h^{3/2}}{6}\mathbb{E}(\text{d}u(D_{V}^{2}V))+\frac{h^{3/2}}{2}\mathbb{E}(D^{2}u(V,D_{V}V)).$
We note that the remainder terms include a linear combination of derivatives of $u$ which can be estimated using (2.14) together with boundedness of $\nabla\phi$ thanks to compactness of $M$ and Assumptions Assumption and Assumption. Due to Assumption Assumption and again using Assumptions  Assumption and Assumption, the last two terms are also bounded by $Ce^{-\lambda(T-t)}h^{2}$, where $C$ and $\lambda$ are independent of $h$ and $T$. Consequently, we arrive at (C.2). ∎
@@
(C.15) $\gamma_{V_{n}}(0)$ $=X_{n}^{h},$
(C.16) $\dot{\gamma}_{V_{n}}(0)$ $=V_{n}:=-\frac{h^{1/2}}{2}\nabla\phi_{n}+g^{-1/2}_{n}\xi_{n+1}.$
Applying the covariant Taylor expansion (B) with $s=\sqrt{h}$ to $u_{n+1}$ around $X_{n}^{h}$, we obtain
@@
(C.17) $+\frac{h^{2}}{24}D^{4}u(\gamma_{V_{n}}(\alpha))(\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha)),$
@@
(C.18) $+\frac{h^{2}}{24}\mathbb{E}[D^{4}u(\gamma_{V_{n}}(\alpha))(\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha))\;|\;X_{n}^{h}],$
where $\partial_{i}$, $i=1,\dots,q$, denote basis vectors of $T_{X_{n}^{h}}M$ and $V_{n}\in T_{X_{n}^{h}}M$. The arguments employed to obtain (C.18) are based on (C.10)-(C.12) except that here we take expectation conditional on $X_{n}^{h}$ and $u$ represents the solution of the Poisson PDE (2.16). Rearranging (C.18) and using the Poisson PDE (2.16), we arrive at
@@
(C.19) $\quad-\frac{h^{2}}{24}\mathbb{E}[D^{4}u(\gamma_{V_{n}}(\alpha))(\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha))\;|\;X_{n}^{h}].$
@@
To show the bound (4.7), we start with (C.17)
@@
(C.20) $\;\;\;+\frac{h^{3/2}}{6}D^{3}u_{n}(V_{n},V_{n},V_{n})+\frac{h^{2}}{24}D^{4}u(\gamma_{V_{n}}(\alpha))(\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha)).$
We rearrange (C.20), sum over $n=0,\dots,N-1$ and use (2.16) to arrive at
@@
(C.21) $\quad-\frac{h^{2}}{24}D^{4}u(\gamma_{V_{n}}(\alpha))(\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha))\bigg{)}.$
@@
(C.22) $\;\;\;+\frac{CNh^{4}}{T^{2}}\sum\limits_{n=0}^{N-1}\big{(}D^{4}u(\gamma_{V_{n}}(\alpha))(\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha))\big{)}^{2}.$
Using (C.8), we get
@@
(C.24) $\mathbb{E}\bigg{(}\sum\limits_{n=0}^{N-1}D^{2}u_{n}(g^{-1/2}_{n}\xi_{n+1},\nabla\phi_{n})\bigg{)}^{2}=\sum\limits_{n=0}^{N-1}\mathbb{E}(D^{2}u_{n}(g^{-1/2}_{n}\xi_{n+1},\nabla\phi_{n}))^{2}.$
@@
(C.25) $=\sum\limits_{n=0}^{N-1}\mathbb{E}(D^{3}u_{n}(\partial_{i},\partial_{j},\partial_{k})(g^{-1/2}_{n}\xi_{n+1})^{i}(g^{-1/2}_{n}\xi_{n+1})^{j}(g^{-1/2}_{n}\xi_{n+1})^{k})^{2}$
@@
(C.26) $=\sum\limits_{n=0}^{N-1}\mathbb{E}(D^{2}u_{n}(g^{-1/2}_{n}\xi_{n+1},g^{-1/2}_{n}\xi_{n+1})-\Delta_{M}u_{n})^{2}.$
Taking expectation on both sides of (C.22), using (C.23)-(C.26) and applying (2.17), we obtain

### v3

@@ 6.2. Sampling on the manifold of SPD matrices
We consider the non-compact manifold of symmetric positive definite (SPD) matrices, and demonstrate numerically that the proposed Riemannian Langevin algorithm results in the desired error bounds when sampling from two distributions which satisfy the assumptions discussed in Section 5. In contrast to the unit sphere in Section 6.1 , the algorithm is implemented using a closed form expression for the exponential map.
Consider the general linear group $GL(m)$ consisting of the set of $m\times m$ real invertible matrices equipped with the group operation of matrix multiplication. Denote by $M=\mathcal{P}_{m}$, the manifold of $m\times m$ symmetric positive definite matrices $\mathcal{P}_{m}=\{X\in GL(m):X^{\intercal}=X,\;y^{\intercal}Xy>0,\;y\in\mathbb{R}^{m}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\setminus\{0\}}\}$, an open cone of dimension $q=m(m+1)/2$.
@@
Affine invariance of the metric ensures that the Riemannian distance $\rho$ and volume form $\mathrm{dvol}_{g}$ will also be invariant under affine transformations. The metric $g$ gives $\mathcal{P}_{m}$ a Riemannian structure with non-positive sectional curvature, and $\mathcal{P}_{m}$ is a non-compact, geodesically complete metric space by the Hopf-Rinow theorem (e.g., Jost, 2008). Moreover, by the Cartan-Hadamard theorem (e.g., Kobayashi and Nomizu, 1969), the inverse exponential map is globally defined, and the cut locus of every point is empty, i.e., every two points in $\mathcal{P}_{m}$ have a unique geodesic connecting them. This means that only a single chart is needed to cover $\mathcal{P}_{m}$.
@@
$\rho(X_{1},X_{2})=\sqrt{\sum_{i=1}^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}m}(\log r_{i})^{2}},$
@@
6.2.1. Riemannian-Gaussian distribution
The Riemannian-Gaussian distribution can be viewed as the generalisation of the isotropic normal distribution on $\mathbb{R}^{q}$ to $\mathcal{P}_{m}$, where the Euclidean distance is replaced by the Riemannian distance in the probability density function:
@@ 6.2.1. Riemannian-Gaussian distribution
The Riemannian-Gaussian distribution satisfies Assumptions Assumption and Assumption, as seen in Example 5.2 (i). This ensures that the Langevin diffusion converges to its invariant measure exponentially fast. From (6.8) the gradient of $\phi$ is
@@
Algorithm 6.2 Algorithm to sample from Riemannian-Gaussian distribution on $\mathcal{P}_{3}$ with $O=I_{3}$, $\sigma=\frac{1}{\sqrt{2}}$ and $X_{0}=\mathsf{hvec}^{-1}((2,4,2,1,1,0)^{\intercal})$.
@@
Figure 2. Sampling from the Riemannian–-Gaussian distribution on $\mathcal{P}_{3}$: log-log plot of the estimation error err against $h$. Error bars correspond to Monte Carlo error MCerr. The reference red line has gradient 1. The parameters for the Riemannian-Gaussian are as in Table 3.
@@ 6.2.2. Distribution with non-convex potential
with the normalisation constant $Z_{m}$. This distribution is referred to as one with a double-well potential. Similar to the Riemannian-Gaussian distribution, the constant $Z_{m}$ is obtained by evaluating
@@ 7. Discussion and Concluding Remarks
The main focus is on sampling from, and computing expectation of functions of, a target distribution. As a consequence, we demonstrate how the concept of weak approximations for SDEs in $\mathbb{R}^{q}$ and the flat torus $\mathbb{T}^{q}$, and the corresponding proof techniques in deriving weak error bounds, can be transferred to the case of compact Riemannian manifolds via the covariant Taylor expansion. The PDE-based proof technique enables obtaining weak error bounds that match the optimal bounds in $\mathbb{R}^{q}$ and flat manifolds; these are superior to the weak error bound currently available in the literature, which follow by first deriving strong approximation (e.g., convergence in expected squared-distance) error bounds. The price we pay for our approach is in the absence of explicit dimension-dependent constants in the derived bounds. The Talay-Tubaro extrapolation in Section 4.4 may be used to extract more information on the constants.
@@ Acknowledgements
The authors were supported by EPSRC grant no. EP/X022617/1. KB acknowledges support from grants EPSRC EP/V048104/1, NSF 2015374 and NIH R37-CA21495. AS acknowledges support by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
Appendix A
@@
(A.1) $F(s)=F(0)+s\frac{\text{d}}{\text{d}s}F(0)+\frac{s^{2}}{2}\frac{\text{d}^{2}}{\text{d}s^{2}}F(0)+\frac{s^{3}}{6}\frac{\text{d}^{3}}{\text{d}s^{3}}F(0)+\frac{s^{4}}{24}\frac{\text{d}^{4}}{\text{d}s^{4}}F(\alpha),$
@@
(A.2) $\dfrac{\text{d}}{\text{d}s}F(s)=\dfrac{\text{d}}{\text{d}s}f(c(s))=\text{d}f(c(s))(\dot{c}(s)).$
@@
(A.4) $\;\;\;+\text{d}f(c(s))(D_{\dot{c}(s)}D_{\dot{c}(s)}\dot{c}(s)),$
where we have utilized the fact that second covariant derivative of functions is symmetric $(0,2)-$tensor. Evaluating (A.2), (A), (A) at $s=0$ and substituting them into (A.1), we obtain the following Taylor formula of $f$ along the curve $c$:
@@
(A.5) $\;\;\;+\dfrac{s^{4}}{24}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dfrac{\text{d}^{4}}{\text{d}s^{4}}f(c(\alpha))},$
@@
Appendix B Proofs of Theorems
@@
Lemma B.1 (One-step error).
@@
(B.3) $u(t+h,c(s))=u(t,c(s))+h\frac{\partial u}{\partial t}(t,c(s))+\frac{h^{2}}{2}\frac{\partial^{2}u}{\partial t^{2}}(t(\alpha_{1}),c(s)),$
@@
Applying the covariant Taylor formula (A) to the function $u(t,c(s))$ along the curve $c$, we get
@@
(B.4) $+\dfrac{s^{4}}{24}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dfrac{\text{d}^{4}}{\text{d}s^{4}}u(t,c(\alpha_{2}))},$
@@
(B.5) $\;\;\;+\frac{s^{2}}{2}\bigg{(}D^{2}\frac{\partial u}{\partial t}(t,c(\alpha_{3}))(\dot{c}(\alpha_{3}),\dot{c}(\alpha_{3}))+\text{d}\dfrac{\partial u}{\partial t}(t,c(\alpha_{3}))(D_{\dot{c}(\alpha_{3})}\dot{c}(\alpha_{3}))\bigg{)},$
@@
$+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dfrac{s^{4}}{24}\dfrac{\text{d}^{4}}{\text{d}s^{4}}u(t,c(\alpha_{2}))}$
@@
(B.6) $+\frac{h^{2}}{2}\frac{\partial^{2}u}{\partial t^{2}}(t(\alpha_{1}),c(s)).$
@@
$+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dfrac{h^{2}}{24}\dfrac{\text{d}^{4}}{\text{d}s^{4}}u(t,c(\alpha_{2}))}$
@@
(B.7) $+\frac{h^{2}}{2}\frac{\partial^{2}u}{\partial t^{2}}(t(\alpha_{1}),c(s))\bigg{]}.$
@@
(B.8) $\mathbb{E}[\xi^{i}]=0,\;\mathbb{E}[\xi^{i}\xi^{j}]=\delta_{ij},\;\mathbb{E}[\xi^{i}\xi^{j}\xi^{k}]=0,\;\mathbb{E}[(\xi^{i})^{2}(\xi^{j})^{2}]=1\text{ for }i\neq j,$
@@
(B.9) $=-\frac{h^{3/2}}{8}(\nabla\phi)^{i}(\nabla\phi)^{j}(\nabla\phi)^{k}-\frac{3h^{1/2}}{2}(\nabla\phi)^{i}g^{jk},$
@@
(B.10) $\mathbb{E}(h^{1/2}Du(V))$ $=-\frac{h}{2}g(\nabla\phi,\nabla u),$
@@
(B.11) $=h\Delta_{M}u+\frac{h^{2}}{4}D^{2}u(\nabla\phi,\nabla\phi),$
@@
(B.12) $\;\;\;\;-\frac{3h^{2}}{2}D^{3}u(\partial_{i},\partial_{j},\partial_{k})(\nabla\phi)^{i}g^{jk},$
(B.13) $\mathbb{E}\Big{(}h^{3/2}D\frac{\partial u}{\partial t}(V)\Big{)}$ $=h^{3/2}D\frac{\partial u}{\partial t}(\partial_{i})\mathbb{E}(V^{i})=\frac{h^{2}}{2}D\frac{\partial u}{\partial t}(\partial_{i})(\nabla\phi)^{i}.$
Substituting (B.10)-(B.13) in (B.7), we get
@@
$\;\;\;\;+\mathbb{E}\bigg{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dfrac{h^{2}}{24}\dfrac{\text{d}^{4}}{\text{d}s^{4}}u(t,c(\alpha_{2}))}$
@@
(B.14) $\;\;\;\;+\frac{h^{3/2}}{6}\mathbb{E}(\text{d}u(D_{V}^{2}V))+\frac{h^{3/2}}{2}\mathbb{E}(D^{2}u(V,D_{V}V)).$
We note that the remainder terms include a linear combination of derivatives of $u$ which can be estimated using (2.14) together with boundedness of $\nabla\phi$ thanks to compactness of $M$ and Assumptions Assumption and Assumption. Due to Assumption Assumption and again using Assumptions  Assumption and Assumption, the last three terms are also bounded by $Ce^{-\lambda(T-t)}h^{2}$, where $C$ and $\lambda$ are independent of $h$ and $T$. Consequently, we arrive at (B.2). ∎
@@
(B.15) $\gamma_{V_{n}}(0)$ $=X_{n}^{h},$
(B.16) $\dot{\gamma}_{V_{n}}(0)$ $=V_{n}:=-\frac{h^{1/2}}{2}\nabla\phi_{n}+g^{-1/2}_{n}\xi_{n+1}.$
Applying the covariant Taylor expansion (A) with $s=\sqrt{h}$ to $u_{n+1}$ around $X_{n}^{h}$, we obtain
@@
(B.17) $+\frac{h^{2}}{24}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dfrac{\text{d}^{4}}{\text{d}s^{4}}u(\gamma_{V_{n}}(\alpha))},$
@@
(B.18) $+\frac{h^{2}}{24}\mathbb{E}[D^{4}u(\gamma_{V_{n}}(\alpha))(\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha))\;|\;X_{n}^{h}],$
where $\partial_{i}$, $i=1,\dots,q$, denote basis vectors of $T_{X_{n}^{h}}M$ and $V_{n}\in T_{X_{n}^{h}}M$. The arguments employed to obtain (B.18) are based on (B.10)-(B.12) except that here we take expectation conditional on $X_{n}^{h}$ and $u$ represents the solution of the Poisson PDE (2.16). Rearranging (B.18) and using the Poisson PDE (2.16), we arrive at
@@
(B.19) $\quad-\frac{h^{2}}{24}\mathbb{E}[D^{4}u(\gamma_{V_{n}}(\alpha))(\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha))\;|\;X_{n}^{h}].$
@@
To show the bound (4.7), we start with (B.17)
@@
(B.20) $\;\;\;+\frac{h^{3/2}}{6}D^{3}u_{n}(V_{n},V_{n},V_{n})+\frac{h^{2}}{24}D^{4}u(\gamma_{V_{n}}(\alpha))(\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha)).$
We rearrange (B.20), sum over $n=0,\dots,N-1$ and use (2.16) to arrive at
@@
(B.21) $\quad-\frac{h^{2}}{24}D^{4}u(\gamma_{V_{n}}(\alpha))(\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha))\bigg{)}.$
@@
(B.22) $\;\;\;+\frac{CNh^{4}}{T^{2}}\sum\limits_{n=0}^{N-1}\big{(}D^{4}u(\gamma_{V_{n}}(\alpha))(\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha),\dot{\gamma}_{V_{n}}(\alpha))\big{)}^{2}.$
Using (B.8), we get
@@
(B.24) $\mathbb{E}\bigg{(}\sum\limits_{n=0}^{N-1}D^{2}u_{n}(g^{-1/2}_{n}\xi_{n+1},\nabla\phi_{n})\bigg{)}^{2}=\sum\limits_{n=0}^{N-1}\mathbb{E}(D^{2}u_{n}(g^{-1/2}_{n}\xi_{n+1},\nabla\phi_{n}))^{2}.$
@@
(B.25) $=\sum\limits_{n=0}^{N-1}\mathbb{E}(D^{3}u_{n}(\partial_{i},\partial_{j},\partial_{k})(g^{-1/2}_{n}\xi_{n+1})^{i}(g^{-1/2}_{n}\xi_{n+1})^{j}(g^{-1/2}_{n}\xi_{n+1})^{k})^{2}$
@@
(B.26) $=\sum\limits_{n=0}^{N-1}\mathbb{E}(D^{2}u_{n}(g^{-1/2}_{n}\xi_{n+1},g^{-1/2}_{n}\xi_{n+1})-\Delta_{M}u_{n})^{2}.$
Taking expectation on both sides of (B.22), using (B.23)-(B.26) and applying (2.17), we obtain